# Parameter Estimation for Weakly Interacting Hypoelliptic Diffusions

### v1

@@ B.2 Proof of Lemma 2
Indeed, we have from Lemmas 1 and 2 that, for any $\theta\in\Theta$:
@@
where $e_{j-1}^{(\iota),k_{1}}\in\mathcal{S}_{t_{j-1}},\,\iota=1,2$. We next prove (19) for any fixed $\theta$. We write the LHS of (19) as $\textstyle\sum_{j=1}^{n}H_{j}^{N,k_{1}k_{2}}(\theta)$ with $\textstyle H_{j}^{N,k_{1}k_{2}}(\theta):=\tfrac{\Delta_{n}}{N}\sum_{i=1}^{N}g(Z_{j-1}^{[i],\theta})\,\mathbf{m}_{j-1}^{[i],k_{1}}\,\mathbf{m}_{j-1}^{[i],k_{2}}.$ Then, Lemmas 1-2 give that:
@@ Appendix C Proof of Proposition 1
where $h$ is a non-decreasing function s.t. $|h(x)|\lesssim x,\,x>0$. The last three terms in the above expression converge to $0$ in probability due to Lemmas 1 and 2. We then conclude from Lemmas 1 and 2, the consistency of $\hat{\theta}^{n,N}$ by noticing that the terms involving $b_{S}$ also converge to $0$ when $\theta=\hat{\theta}^{n,N}$.
@@
We then conclude from Lemmas 1, 2, and the consistency of the estimator.
@@
for some $f_{\ell_{1},\ell_{2}}^{k_{1},k_{2}},\,f_{\ell}^{k_{1},k_{2}}\in\mathcal{R}$. From Lemmas 1, 2 we have that:
@@ D.1 Proof of (57)
From Lemma 1 and (36) in Lemma 2 we obtain that $\textstyle\mathbb{E}_{\theta^{\dagger}}\bigl{|}\mathscr{J}_{n,N}^{k_{1},(1)}\bigr{|}\leq C\sqrt{\tfrac{\Delta_{n}}{N}}\sum_{i=1}^{N}\sum_{j=1}^{n}\Delta_{n}^{3/2}=C\times T\sqrt{N\Delta_{n}^{2}},$ for some constant $C>0$ depending on $T>0$ but not on $n,N$. We also have from Lemma 1 and (37) in Lemma 2 that:
@@ D.2 Proof of (58)
for some $e_{j-1}^{[i_{1},i_{2}]}\in\mathcal{S}_{t_{j-1}}$, where in the second line, we have used equation (37) from Lemma 2 and in the third line, we have used the equality:
@@
Case II. $(k_{1},k_{2})\in\mathcal{I}_{\beta}\times\mathcal{I}_{\beta}$. We have from Lemma 2 that $\textstyle\mathscr{J}_{n,N}^{k_{1},k_{2}}=\sum_{\iota=1,2,3}G_{k_{1},k_{2}}^{(\iota)}$ with:
@@
We immediately see that $G_{k_{1},k_{2}}^{(2)}$ converges to $0$ as $N\Delta_{n}\to 0$ in $L_{1}$. We then study the first term. We apply (37) and (38) in Lemma 2 to get that:
@@
Case III. $(k_{1},k_{2})\in\mathcal{I}_{\alpha_{S}}\times\mathcal{I}_{\alpha_{S}}$ . Making use of Lemma 2 and the estimate (62), we get $\textstyle\mathscr{J}_{n,N}^{k_{1},k_{2}}=\sum_{\iota=1,2}H_{k_{1},k_{2}}^{(\iota)}$ with:
@@
We have from Lemma 2 that:
@@
It follows from Lemma 2 that:
@@ D.3 Proof of (59)
for some $e_{j-1}^{[i_{1},i_{2},i_{3},i_{4}]}\in\mathcal{S}_{t_{j-1}}$. We immediately have that $\widetilde{\mathscr{J}}_{n,N}^{\,k_{1},(3)}$ converges to $0$ as $N\Delta_{n}\to 0$ under the $L_{1}$-norm. For the first term, we apply (38) in Lemma 2 to get that:
@@
Due to Lemma 1, we immediately obtain that $\widetilde{\mathscr{J}}_{n,N}^{k_{1},(1)}\to 0$ under $L_{1}$ as $N\Delta_{n}\to 0$. We thus conclude proof of Case III., and now the proof of Proposition 2 is complete.
Proof of Lemma 1. Using the decomposition $\mathbf{m}_{j-1}^{[i]}=C_{j-1}^{[i]}+D_{j-1}^{[i]}$ specified in (40) and noticing that $\mathbb{E}_{\theta^{\dagger}}|D_{j-1}^{[i]}|^{p}=\mathcal{O}(\Delta_{n}^{1/2}),\,p\geq 1$, we have that (LHS of (LABEL:eq:prod_four_beta)) $=\textstyle\sum_{\iota=1,2}E^{\gamma,(\iota)}_{j-1}+g_{j-1}^{\gamma}(\Delta_{n})$ for some $g_{j-1}^{\gamma}\in\mathcal{S}_{t_{j-1}}$, where:

### v2

@@ B.2 Proof of Lemma 3
Indeed, we have from Lemmas 2 and A.2 that, for any $\theta\in\Theta$:
@@
where $e_{j-1}^{(\iota),k_{1}}\in\mathcal{S}_{t_{j-1}},\,\iota=1,2$. We next prove (20) for any fixed $\theta$. We write the LHS of (20) as $\textstyle\sum_{j=1}^{n}H_{j}^{N,k_{1}k_{2}}(\theta)$ with $\textstyle H_{j}^{N,k_{1}k_{2}}(\theta):=\tfrac{\Delta_{n}}{N}\sum_{i=1}^{N}g(Z_{j-1}^{[i],\theta})\,\mathbf{m}_{j-1}^{[i],k_{1}}\,\mathbf{m}_{j-1}^{[i],k_{2}}.$ Then, Lemmas 2-A.2 give that:
@@ Appendix C Proof of Proposition 1
where $h$ is a non-decreasing function s.t. $|h(x)|\lesssim x,\,x>0$. The last three terms in the above expression converge to $0$ in probability due to Lemmas 2 and 3. We then conclude from Lemmas 2 and 3, the consistency of $\hat{\theta}^{n,N}$ by noticing that the terms involving $b_{S}$ also converge to $0$ when $\theta=\hat{\theta}^{n,N}$.
@@
We then conclude from Lemmas 2, 3, and the consistency of the estimator.
@@
for some $f_{\ell_{1},\ell_{2}}^{k_{1},k_{2}},\,f_{\ell}^{k_{1},k_{2}}\in\mathcal{R}$. From Lemmas 2, 3 we have that:
@@ D.1 Proof of (58)
From Lemma A.1 and (37) in Lemma A.2 we obtain that $\textstyle\mathbb{E}_{\theta^{\dagger}}\bigl|\mathscr{J}_{n,N}^{k_{1},(1)}\bigr|\leq C\sqrt{\tfrac{\Delta_{n}}{N}}\sum_{i=1}^{N}\sum_{j=1}^{n}\Delta_{n}^{3/2}=C\times T\sqrt{N\Delta_{n}^{2}},$ for some constant $C>0$ depending on $T>0$ but not on $n,N$. We also have from Lemma A.1 and (38) in Lemma A.2 that:
@@ D.2 Proof of (59)
for some $e_{j-1}^{[i_{1},i_{2}]}\in\mathcal{S}_{t_{j-1}}$, where in the second line, we have used equation (38) from Lemma A.2 and in the third line, we have used the equality:
@@
Case II. $(k_{1},k_{2})\in\mathcal{I}_{\beta}\times\mathcal{I}_{\beta}$. We have from Lemma A.2 that $\textstyle\mathscr{J}_{n,N}^{k_{1},k_{2}}=\sum_{\iota=1,2,3}G_{k_{1},k_{2}}^{(\iota)}$ with:
@@
We immediately see that $G_{k_{1},k_{2}}^{(2)}$ converges to $0$ as $N\Delta_{n}\to 0$ in $L_{1}$. We then study the first term. We apply (38) and (39) in Lemma A.2 to get that:
@@
Case III. $(k_{1},k_{2})\in\mathcal{I}_{\alpha_{S}}\times\mathcal{I}_{\alpha_{S}}$ . Making use of Lemma A.2 and the estimate (63), we get $\textstyle\mathscr{J}_{n,N}^{k_{1},k_{2}}=\sum_{\iota=1,2}H_{k_{1},k_{2}}^{(\iota)}$ with:
@@
We have from Lemma A.2 that:
@@
It follows from Lemma A.2 that:
@@ D.3 Proof of (60)
for some $e_{j-1}^{[i_{1},i_{2},i_{3},i_{4}]}\in\mathcal{S}_{t_{j-1}}$. We immediately have that $\widetilde{\mathscr{J}}_{n,N}^{\,k_{1},(3)}$ converges to $0$ as $N\Delta_{n}\to 0$ under the $L_{1}$-norm. For the first term, we apply (39) in Lemma A.2 to get that:
@@
Due to Lemma D.1, we immediately obtain that $\widetilde{\mathscr{J}}_{n,N}^{k_{1},(1)}\to 0$ under $L_{1}$ as $N\Delta_{n}\to 0$. We thus conclude proof of Case III., and now the proof of Proposition 2 is complete.
Proof of Lemma D.1. Using the decomposition $\mathbf{m}_{j-1}^{[i]}=C_{j-1}^{[i]}+D_{j-1}^{[i]}$ specified in (41) and noticing that $\mathbb{E}_{\theta^{\dagger}}|D_{j-1}^{[i]}|^{p}=\mathcal{O}(\Delta_{n}^{1/2}),\,p\geq 1$, we have that (LHS of (LABEL:eq:prod_four_beta)) $=\textstyle\sum_{\iota=1,2}E^{\gamma,(\iota)}_{j-1}+g_{j-1}^{\gamma}(\Delta_{n})$ for some $g_{j-1}^{\gamma}\in\mathcal{S}_{t_{j-1}}$, where: